OpenStudy (anonymous):

limite x-> ∞ x/sqrt x+1

OpenStudy (anonymous):

Do you know what the conjugate is?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ x }{ \sqrt{1+x} }\]

OpenStudy (anonymous):

no raiz de x+1

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \sqrt{\frac{ x^2 }{ 1+x }}\] and then you can use power law, and L'Hospital's rule.

OpenStudy (anonymous):

\[limite \rightarrow \infty \]

OpenStudy (anonymous):

\[\frac{ x }{ \sqrt{x+1} }\]

OpenStudy (anonymous):

understand ???

OpenStudy (anonymous):

k

OpenStudy (anonymous):

can you help me ?

OpenStudy (anonymous):

heeelp